OpenStudy (anonymous):

If George drives at a steady rate of 65 miles per hour, how many miles will he travel in 1.5 hours?

OpenStudy (anonymous):

Well in one hour he will obviously be driving 65 miles. So, 65 right there. Then to find how much he drove in the remainding half an hour, just cut 65 in half. 65/2 = 32.5 65 + 32.5 = 97.5 Or you could just 65 * 1.5 = 97.5

OpenStudy (cggurumanjunath):

s=d/t d=s*t=65*1.5

OpenStudy (cggurumanjunath):

65*1.5 miles